# Graphics Processing Unit DGF Dense Geometry Format

### v1

@@ 4 Results and Discussion
We implemented the proposed method in DirectX 12 and HLSL. Rendering each frame consists of the following five passes: (1) G-buffer generation, (2) training data generation, (3) network training, (4) inference & light sampling, and (5) shading. The neural network is implemented as a fully-fused MLP [MRNK21], allowing us to train and infer the neural network entirely in the on-chip shared memory. Our implementation allows to execute inference inline in the scope of DXR ray tracing shaders, enabling to use our method on every bounce of light while tracing paths. All tests were executed on AMD Radeon RX 7900 XT.
@@
Figure 5: A comparison of NRC (top), our neural light sampling (NLS) (center) and screen space ReSTIR (bottom) on a progression of 128 frames. Number under each image represents its FLIP error, which decreases over time for all methods, but NRC suffers from color artifacts, especially at the beginning of training. Our method remains unbiased even at the beginning while the network is not sufficiently trained, exhibiting as increased variance, similarly to ReSTIR. Our method reduces the FLIP error faster than other methods, achieving lower error than ReSTIR even only after three training steps
@@ 4.1 Comparison to Screen-space ReSTIR
As a reference method, we implemented screen-space ReSTIR using 8 initial candidates, casting a shadow ray for the selected initial candidate. We use temporal resampling where we clamp contribution of the previous reservoir to 20×20$\times$20 × the contribution of new reservoir. Spatial resampling uses a radius of 32 pixels. Unless stated otherwise, we use one sample per pixel for all tests, and both ReSTIR and the neural network are converged for 1024 frames. Note that the FLIP error of both methods is already significantly reduced after 20 to 30 training steps (see Figure 5).
Compared to screen-space ReSTIR, our neural light sampling achieves a lower FLIP error [ANA21] at a similar time budget, especially in the occluded regions (see Figures Neural Visibility Cache for Real-Time Light Sampling and 8). FLIP error is reduced by about $20\%$ for the Kitchen scene and about $45\%$ for the Sponza scene. We can make the method unbiased by clamping the output of the neural network (discussed in Section 3.1) at the cost of slightly increased variance (see Figure 4). Neural DI that directly uses the visibility estimates provided by the neural network (see Section 3.1.1) is biased by definition, yet it provides significantly lower error than ReSTIR and neural light sampling (see Figure Neural Visibility Cache for Real-Time Light Sampling). The bias exhibits at the edges and heavily shadowed areas (see Figure 2).
4.1.1 Combining ReSTIR with Clustered NVC
@@
Figure 6: A comparison of standard ReSTIR (left), ReSTIR with initial candidates generated by our clustered NVC (center) and ground truth (right) in an idealized case when disocclusions happen in every pixel. The Subway scene contains $60k$ lights. We simulate disocclusions by invalidating motion vectors for all pixels, using one sample per pixel. This demonstrates how clustered NVC can help ReSTIR to recover after disocclusion.
To generate initial candidates, standard ReSTIR uses a RIS loop with user-selected number of candidates (we use $8$), which generates a reservoir that can participate in the streaming RIS and for spatial and temporal reuse. We combine our neural approach with ReSTIR by replacing the initial candidates generation routine with our clustered NVC. Our method also generates a reservoir, using a two-step process described in section 3.4, therefore it can be directly used to generate initial candidates.
While using NVC for initial candidates improves the quality of ReSTIR overall, the biggest benefit is for boosting convergence rate for disocclusions. Figure 6 shows a significant noise reduction when NVC is used for initial candidates after a disocclusion event. Running NVC for ReSTIR once it converges only improves ReSTIR quality insignificantly, while introducing an overhead of running the inference for every pixel. Therefore, we recommend running NVC only for disoccluded pixels to boost quality of initial candidates. This implementation leads to runtime performance only $5\%$ lower compared to standard ReSTIR, on a test using a dynamic camera flying around the scene for 1024 frames.
@@ 4.2 Comparison to Neural Radiance Cache
Figure 7: NRC approximating a product of radiance and visibility exhibits blurry artifacts and color shifts (left). Our method computes exact radiance analytically and samples the light sources stochastically for unbiased rendering (middle). The ground truth accumulated with many frames (right).
For a direct comparison of our method to the neural radiance cache (NRC) [MRNK21], we have implemented a version of NRC which caches only direct illumination from all lights on a primary hit (NRC DI). We use the exactly same network architecture (except the output layer which consists of three neurons as radiance is RGB value, and we use leaky ReLU activation function to allow for unbounded values) and training data procedure as we do for our method. The NRC is not limited by number of lights, but it has several drawbacks compared to our method. For optimal results, NRC implementation requires a larger MLP, which lowers the training and inference performance. Another disadvantage is that NRC requires more training steps to achieve usable results (see Figure 5). At the beginning of training, the NRC produces random colors, introducing significant error to the resulting image. NRC predicts a product of the wavelength-dependent radiance and visibility, manifesting blurry artifacts and color shifts (see Figure 7). Our method instead calculates the radiance analytically with LTCs. The predicted visibility is used to guide the light sampling process, which remains unbiased, even when the network training has not yet converged (in this case, we get increase in variance instead of bias). Compared to NRC, our method is limited to direct illumination and a smaller number of lights, but for this purpose it achieves unbiased results with a smaller network and does not suffer from artifacts due to under-training and radiance approximation.
@@ 4.3 Performance
Training data generation with our default configuration (8196 training examples and 32 lights) takes 0.34 ms for the Kitchen scene and 0.50 ms for the Sponza scene. The neural network training step (backpropagation and optimization) takes 0.75 ms for both scenes. When training data generation and training itself is implemented as separate steps, where first step writes out the data into memory for second step to consume, it is possible to perform multiple training iterations over same data. As we only perform one step per frame, we can fuse data generation and training together, achieving about $5\%$ speedup. We run inference once per pixel at 1.87 ms per frame for $1920\times 1080$ resolution with 32 lights. The inference itself takes 1.32 ms, while the remaining time is spent on the WRS algorithm. For comparison, replacing the inference call of our NVC by casting a shadow ray towards each of the light sources for every pixel on screen takes about $7$ ms for the Kitchen scene with 32 lights (more than 3×3$\times$3 × higher).
@@ 5 Conclusion and Future Work
We proposed a lightweight neural-based sampling method for real-time direct illumination based on caching the non-binary visibility. With a minor modification, our method provides unbiased estimates with lower error than screen-space ReSTIR at a similar cost. Compared to ReSTIR, our method operates in the world-space, making it more robust to visibility changes that ReSTIR struggles with. In fact, our neural light sampling could be used in combination with ReSTIR to sample the initial candidates. As a world-space method, it can also be used for direct illumination of volumes of participating media. We used a clustered approach to support an arbitrary number of lights with a fixed-size neural network. We also proposed a biased variant that directly uses the visibility estimates, decreasing variance even further at the cost of some bias. Thanks to continuous online training, our method adapts to the dynamic content including animated lights, geometry, and camera. Compared to neural importance sampling of many lights [FHBK25], we are only caching visibility with a smaller MLP, achieving faster training and simpler implementation (except from the position, we do not require other inputs, e.g. the surface normal to the MLP).
The limitation of our clustered NVC approach is that efficiency is highly dependent on the quality of the clusters created, and total number of lights. For future work, an interesting direction would be to explore other methods for light clustering, e.g., the ones based on hierarchies [VKK18, FHBK25] and light culling [TH16]. An early approach of Shirley et al. [SWZ96] of sorting lights into sets of important and unimportant lights could be adapted to our approach by creating one cluster of unimportant lights to ensure unbiasedness and use remaining available clusters to represent the important lights.
Another interesting research direction would be to cache mutual visibility between any two points in the scene. This would enable us to query visibility of any light, not just the ones represented by output neurons. To make this practical, such query would either have to be faster than a ray cast, or it would have to return visibility for a batch of queries in a single inference call, to amortize the cost.
Finally, we need to overcome the limitation of the count of lights (or light clusters) due to the restrictions imposed on the size of the neural network. We can achieve linear scaling naïvely by having multiple networks and possibly time-slicing their training (only train one network per frame) to maintain fixed training budget. Overcoming this limitation in a scalable way is a non-trivial task for the future work.
@@ Acknowledgements
We would like to thank Holger Grün and Carsten Benthin for useful discussions and their support. Model courtesy: Sponza (Crytek), Subway (Sketchfab/Alex Murias), Country Kitchen (Blendswap/Jay-Artist).

### v2

@@ 4 Results and Discussion
We implemented the proposed method in DirectX 12 and HLSL. Rendering each frame consists of the following five passes: (1) G-buffer generation, (2) training data generation, (3) network training, (4) inference and light sampling, and (5) shading. The neural network is implemented as a fully fused MLP Müller et al. (2021), allowing us to train and infer the neural network entirely in the on-chip shared memory. Our implementation allows to execute inference inline in the scope of DirectX Raytracing (DXR) ray tracing shaders, enabling to use our method on every bounce of light while tracing paths. All tests were executed on AMD Radeon RX 7900 XT.
@@
Refer to caption Refer to caption
Ground Truth FLIP Over Time
Figure 5: A comparison of NRC (top), our neural light sampling (NLS) (middle) and screen space ReSTIR (bottom) on a progression of 128 frames. The number under each image represents its FLIP error, which decreases over time for all methods, but NRC suffers from color artifacts, especially at the beginning of training. Our method remains unbiased even at the beginning while the network is not sufficiently trained, exhibiting as increased variance, similarly to ReSTIR. Our method reduces the FLIP error faster than other methods, achieving lower error than ReSTIR even only after three training steps.
@@ 4.1 Comparison to Screen-Space ReSTIR
As a reference method, we implemented screen-space ReSTIR using eight initial candidates, casting a shadow ray for the selected initial candidate. We use temporal resampling where we clamp the contribution of the previous reservoir to $20\times$ the contribution of the new reservoir. Spatial resampling uses a radius of 32 pixels. Unless stated otherwise, we use one sample per pixel for all tests, and both ReSTIR and the neural network are converged for 1024 frames. Note that the FLIP error of both methods is already significantly reduced after 20 to 30 frames (see Figure 5).
Compared to screen-space ReSTIR, our neural light sampling achieves a lower FLIP error Andersson et al. (2021) at a similar time budget, especially in the occluded regions (see Figures Neural Visibility Cache for Real-Time Light Sampling and 6). FLIP error is reduced by about $20\%$ for the Kitchen scene and about $45\%$ for the Sponza scene. We can make the method unbiased by clamping the output of the neural network (discussed in Section 3.1) at the cost of slightly increased variance (see Figure 2). Neural DI that directly uses the visibility estimates provided by the neural network (see Section 3.2) is biased by definition, yet it provides significantly lower error than ReSTIR and neural light sampling (see Figure Neural Visibility Cache for Real-Time Light Sampling). The bias exhibits at the edges and in heavily shadowed areas (see Figure 3).
4.2 Combining ReSTIR with Clustered NVC
@@
Figure 7: A comparison of standard ReSTIR (left), ReSTIR with initial candidates generated by our clustered NVC (center), and the ground truth (right) in an idealized case where disocclusions happen in every pixel. The Subway scene contains 59K lights. We simulate disocclusions by invalidating motion vectors for all pixels, using one sample per pixel. This demonstrates how clustered NVC can help ReSTIR to recover after disocclusion.
To generate initial candidates, standard ReSTIR uses a RIS loop with a user-selected number of candidates (we use eight), which generates a reservoir that can participate in the streaming RIS and for spatial and temporal reuse. We combine our neural approach with ReSTIR by replacing the initial candidates generation routine with our clustered NVC. Our method also generates a reservoir, using a two-step process described in Section 3.6, therefore it can be directly used to generate initial candidates.
Though using NVC for initial candidates improves the quality of ReSTIR overall, the biggest benefit is for boosting the convergence rate for disocclusions. Figure 7 shows a significant noise reduction when NVC is used for initial candidates after a disocclusion event. Running NVC for ReSTIR once it converges only improves ReSTIR quality insignificantly; the average FLIP reduction is about $5\%$ (see Figure 8, while introducing an overhead of running the inference for every pixel. Therefore, we recommend running NVC only for disoccluded pixels to boost the quality of initial candidates. This implementation leads to runtime performance being only $5\%$ lower compared to standard ReSTIR, on a test using a dynamic camera flying around the scene for 1024 frames. Results of clustered NVC in comparison to other methods are summarized in Figure 8.
Subway ZeroDay Bistro Exterior
59,164 lights 4820 lights 4499 lights
RIS
Refer to caption Refer to caption Refer to caption
0.880 0.828 0.836
2.92 ms 2.00 ms 2.74 ms
C-NVC (ours)
Refer to caption Refer to caption Refer to caption
0.768 0.754 0.790
7.37 ms 6.18 ms 7.85 ms
ReSTIR
Refer to caption Refer to caption Refer to caption
0.500 0.580 0.445
4.48 ms 3.43 ms 5.15 ms
[…]
Refer to caption Refer to caption Refer to caption
0.464 0.569 0.423
8.99 ms 7.75 ms 8.24 ms
Ground Truth
Refer to caption Refer to caption Refer to caption
Figure 8: Comparison of tested methods on scenes with many lights, captured after 1024 frames. The numbers under each image correspond to the FLIP error against the ground truth and a total frame time. Note that Cluster NVC is more expensive than NVC compared to ReSTIR, due to $6\times$ more training data needed and a more complex sampling process. For scenes with up to 32 lights, the performance of NVC is comparable to ReSTIR. In these tests, we ran Clustered NVC (C-NVC) to generate initial ReSTIR candidates for every pixel, instead of only on disocclusions as we recommend.
@@ 4.3 Comparison to Neural Radiance Cache
For a direct comparison of our method to the neural radiance cache (NRC) Müller et al. (2021), we have implemented a version of NRC that caches only direct illumination from all lights on a primary hit (NRC DI). We use the same network architecture and training data procedure as we do for our method, except the output layer has three neurons for RGB radiance values, and we use the leaky ReLU activation function to allow for unbounded values. NRC is not limited by the number of lights, but it has several drawbacks compared to our method. For optimal results, NRC implementation requires a larger MLP, which increases the training and inference time. Another disadvantage is that NRC requires more training steps to achieve usable results (see Figure 5). At the beginning of training, NRC produces random colors, introducing significant error to the resulting image. NRC predicts a product of the wavelength-dependent radiance and visibility, manifesting blurry artifacts and color shifts (see Figure 9). Our method instead calculates the radiance analytically with LTC. The predicted visibility is used to guide the light sampling process, which remains unbiased, even when the network training has not yet converged. In this case, we get an increase in variance instead of bias. Compared to NRC, our method is limited to direct illumination and a smaller number of lights, but for this purpose it achieves unbiased results with a smaller network and does not suffer from artifacts due to under-training and radiance approximation.
@@
Figure 9: NRC approximating a product of radiance and visibility exhibits blurry artifacts and color shifts (left). Our method computes exact radiance analytically and samples the light sources stochastically for unbiased rendering (middle). The ground truth was accumulated with many frames (right).
@@ 4.4 Performance
The performance of all tested methods is summarized in Table 1. Training data generation with our default configuration (8196 training examples and 32 lights) takes 0.34 ms for the Kitchen scene and 0.43 ms for the Sponza scene. The neural network training step (backpropagation and optimization) takes $\sim 0.75$ ms for both scenes. When training data generation and the training itself are implemented as separate steps, where the first step writes out the data into memory for the second step to consume, it is possible to perform multiple training iterations over the same data. As we only perform one step per frame, we can fuse data generation and training together, achieving about $5\%$ speedup. We run inference once per pixel at $\sim 1.8$ ms per frame for $1920\times 1080$ resolution with 32 lights. The inference itself takes 1.32 ms, while the remaining time is spent on the WRS algorithm. For comparison, replacing the inference call of our NVC by casting a shadow ray toward each of the light sources for every pixel on screen takes about 7 ms for the Kitchen scene with 32 lights (more than $3\times$ higher).
G-Buffer Training (fused) Light Sampling Shading Frame Total
[ms] [ms] [ms] [ms] [ms]
RIS
Subway 1.09 - 0.62 1.16 2.92
ZeroDay 0.61 - 0.36 0.99 2.00
Bistro 0.78 - 0.31 1.61 2.74
Kitchen 0.58 - 0.47 0.57 1.65
Sponza 1.00 - 0.27 0.67 1.97
ReSTIR
Subway 1.17 - 2.01 1.11 4.48
ZeroDay 0.64 - 1.67 0.92 3.43
Bistro 0.83 - 2.30 1.83 5.15
Kitchen 0.58 - 2.45 0.53 3.61
Sponza 1.00 - 2.23 0.64 3.95
NVC (ours)
Kitchen 0.60 0.99 1.74 0.52 3.95
Sponza 0.99 1.18 1.85 0.61 4.69
Clustered NVC (ours)
Subway 1.11 3.01 1.87 1.11 7.37
ZeroDay 0.63 2.63 1.70 0.95 6.18
Bistro 0.80 3.20 1.71 1.87 7.85
[…]
Subway 1.12 3.00 3.61 1.05 8.99
ZeroDay 0.64 2.58 3.43 0.90 7.75
Bistro 0.80 3.38 2.24 1.61 8.24
[…]
@@ 5 Conclusion and Future Work
We proposed a lightweight neural-based sampling method for real-time direct illumination based on caching the nonbinary visibility. With a minor modification, our method provides unbiased estimates with lower error than screen-space ReSTIR at a similar cost. Compared to ReSTIR, our method operates in the world space, making it more robust to visibility changes with which ReSTIR struggles. In fact, our neural light sampling could be used in combination with ReSTIR to sample the initial candidates. As a world-space method, it can also be used for direct illumination of volumes of participating media. We used a clustered approach to support an arbitrary number of lights with a fixed-size neural network. We also proposed a biased variant that directly uses the visibility estimates, decreasing variance even further at the cost of some bias. Thanks to continuous online training, our method adapts to the dynamic content including animated lights, geometry, and camera. Compared to neural importance sampling of many lights Figueiredo et al. (2025), we are only caching visibility with a smaller MLP, achieving faster training and simpler implementation (we do not require other inputs, e.g., the surface normal to the MLP, only the position).
The limitation of our clustered NVC approach is that efficiency is highly dependent on the quality of the clusters created and the total number of lights. For future work, an interesting direction would be to explore other methods for light clustering, e.g., the ones based on hierarchies Vévoda et al. (2018); Figueiredo et al. (2025) and light culling Tokuyoshi and Harada (2016). An early approach of Shirley et al. (1996) of sorting lights into sets of important and unimportant lights could be adapted to our approach by creating one cluster of unimportant lights to ensure unbiasedness and using the remaining available clusters to represent the important lights.
Another interesting research direction would be to cache mutual visibility between any two points in the scene. This would enable us to query the visibility of any light, not just the ones represented by output neurons. To make this practical, such a query would either have to be faster than a ray cast or have to return visibility for a batch of queries in a single inference call, to amortize the cost.
Finally, we need to overcome the limitation of the count of lights (or light clusters) due to the restrictions imposed on the size of the neural network. We can achieve linear scaling naïvely by having multiple networks and possibly time-slicing their training (only train one network per frame) to maintain a fixed training budget. Overcoming this limitation in a scalable way is a nontrivial task for the future.
@@ Acknowledgements
We would like to thank Holger Grün and Carsten Benthin for useful discussions and their support. The models are courtesy of Crytek (Sponza), Sketchfab/Alex Murias (Subway), Blendswap/Jay-Artist (Country Kitchen), Mike Winkelmann (ZeroDay), and Amazon Lumberyard (Bistro).